OpenStudy (anonymous):

7/13 divided by 0 is not possible to divide right. because anything that is divided by 0 is 0??

OpenStudy (anonymous):

You can divide by 0. It's just the rule of maths

OpenStudy (anonymous):

it is infinity not zero

OpenStudy (anonymous):

ok..

OpenStudy (wasiqss):

anything divide by zero is always infinity,

OpenStudy (anonymous):

1/R6?

OpenStudy (anonymous):

division by zero is not defined! when we take the limit making a number to tend to zero, the result will tend to infinite!

OpenStudy (anonymous):

\[\frac{1}{0}=\infty\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0} (a/x) \rightarrow \infty\]

OpenStudy (anonymous):

'a' is nonzero